OpenStudy (anonymous):

if x^2 = 2x + 6 ... i need to solve for x.

OpenStudy (arindameducationusc):

=>X^2-2x-6=0

OpenStudy (anonymous):

You can use the Quadratic Formula to find the values of x

pooja195 (pooja195):

a ,b , c values are \[\huge\rm Ax^2+Bx+C=0\] where a =leading coefficient b= leading coefficient c= constant term \[\huge~\rm~x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (arindameducationusc):

http://www.purplemath.com/modules/quadform.htm

OpenStudy (anonymous):

(x^2-2x+1)=6+1 (x-1)^2=7 now you can solve it.

OpenStudy (arindameducationusc):

then use quadratic formula

OpenStudy (anonymous):

\[x^2 -2x -6 = 0 \] is in the form of \[ax^2 + bx + c = 0\] where a = 1, b = -2, and c = -6

OpenStudy (anonymous):

Simply plug those numbers into the Quadratic Formula

OpenStudy (anonymous):

i think im maybe misunderstanding what im supposed to do? the problem is if x^2= 2x +6 then x can equal (A) 2 - the square root of 28 (B) 1 - the square root of 28 (C) 1 - 2 root 7 (D) 1 - root 7 and none of the above.. i do have to solve for x right? i feel like im missing something